# Plug-and-Play Transformer Modules for Test-Time Adaptation

### 10 Jan 2024 v2

@@ 4.2 Initialization and training of module selector
Drawing from the success of the attention mechanism in achieving impressive few-shot performance in supervised model ensemble scenarios [37, 25], we have extended this observation to test-time adaptation. Even in the unsupervised context of test-time adaptation, we have found that an attention module, which acts as a module selector, can effectively integrate information from multiple source domains into the target domain by dynamically assigning weights and updating on-the-fly. This enables PLUTO to plug modules from the module store and adapt them at test time with few-shot unlabeled data.
@@
In order to convert 2D image input into sequential data for transformers and module selector to process, the image input $x\in\mathbb{R}^{H\times W\times C}$ is reshaped into a sequence of flattened 2D patches [8]$\mathbf{X}_{p}\in\mathbb{R}^{l\times(P^{2}\cdot C)}$, where $(H,W)$ is the height and width of the original image, $C$ is the number of channels, $(P,P)$ is the size of each segmented image patch, and $l=HW/P^{2}$ is number of patches and the effective input sequence length. Assuming that the transformer uses constant latent vector size $d$ through all of its layers, we flatten the sequence of the patches and map them to $d$ dimentions: $\hat{\mathbf{X}}=[\mathbf{X}_{p}\mathbf{E}]_{p=1}^{N},\mathbf{E}\in\mathbb{R}^{(P^{2}\cdot C)\times d}$. The patch embedding projection $\mathbf{E}$ is a (frozen) trainable parameter in the pretrained ViT.
@@
The module selector, which is an attention module, gauges the relevance between the output logits of the source modules and instances $\bm{x}_{i}$ in the current ($t$-th) target (abbreviated as $T$) batch $D_{T}^{(t)}=\{x_{i}^{(t)}\}_{i=1}^{B}\sim\mathcal{D}_{T}^{(t)}$. It then assigns weights based on this relevance. Logits of source models, on the other hand, are a function of both the pretrained module and the pretrained transformer and are also capable of modeling the uncertainty in predictions from different sources. We perform max-pooling over the sequence $\hat{\mathbf{X}}=[\mathbf{X}_{p}\mathbf{E}]_{p=1}^{N}\in\mathbb{R}^{l\times d}$ and obtain $\hat{x}\in\mathbb{R}^{d}$ as the representation of the original image input. Furthermore, we pass the input instance $x$ through $N$ pretrained models and obtain pre-softmax logits $\{l(x_{i})^{j}\}_{j=1}^{N}\in\mathbb{R}^{v}$. With four trainable parameters: $W_{d,x}\in\mathbb{R}^{d\times d^{\prime}_{x}},W_{u,x}\in\mathbb{R}^{d^{\prime}_{x}\times d^{\prime}},W_{d,l}\in\mathbb{R}^{v\times d^{\prime}_{l}},W_{u,l}\in\mathbb{R}^{d^{\prime}_{l}\times d^{\prime}}$, the module selector first project the pre-softmax logits and the representation of the original image input into another representational space: $\mathbf{h}_{x}=W_{u,x}^{\intercal}\cdot\gamma(W_{d,x}^{\intercal}\cdot\hat{x}),\mathbf{h}_{l,j}=W_{u,l}^{\intercal}\cdot\gamma(W_{d,l}^{\intercal}\cdot{l(x)^{j}})$. With application of LN [2] we are able to obtain the final projected representations $\mathbf{h}_{x},\mathbf{h}_{l,j}\in\mathbf{R}^{d^{\prime}}$ and compute the attention weight $w(x)^{j}$ to capture the relationship between the input $x$ and its representation from the pretrained models $l(x)^{j}$:
@@
The ensembled logit is the linear combination of the weighted logits: $l(x)=\sum_{j=1}^{N}w(x)^{j}l(x)^{j}.$ If the ground-truth (GT) label of the target sample $x$ is known (e.g. for source datasets $\{S_{j}\}_{j=1}^{N}$), the module selector $\mathcal{G}$ can be randomly initialized and trained/updated by minimizing its cross-entropy loss between $\text{softmax}(l(x))$ and the GT label $y$. As disucssed in [37], this enables $\mathcal{G}$ to capture the sample-specific preference of different source models.
Otherwise, if the GT label of the target sample $x$ is unknown, we denote its pseudo-label, as predicted by source $j$, as $\hat{y}_{j}=f^{j}_{\theta_{j}}(x)$. We linearly combine these pseudo-labels by attention weights $\{w(x)^{j}\}_{j=1}^{N}$ to get the weighted pseudo-label $\hat{y}=\sum_{j=1}^{N}w(x)^{j}\hat{y}_{j}$. Using these weighted pseudo-labels for all the samples in the current ($t$-th) batch we calculate the expected Shannon entropy as
@@
where $K$ is the number of classes and $\hat{y}_{c}$ is the $c$-th entry of the predicted pseudo-label. By performing pseudo-label entropy minimization on Eqn. 2, we can update the module selector parameters(from $\mathcal{G}$ to $\mathcal{G}^{*}$) and obtain updated attention weights for Eqn. 1.
@@ 4.3 Domain adaptation of LN affine parameters
In TTA, prior methods often conduct adaptation on pre-trained models with batch normalization (BN) layers[22] and most of them are built upon BN statistics adaptation[43, 52, 56]. However, recent works[32, 5, 24] found out that layer normalization (LN, [2]) can be a better adaptation choice for transformers. Specifically, for a layer with $d$-dimensional input $x=(x^{(1)},\cdots,x^{(d)})^{\intercal}\in\mathbb{R}^{d}$, the layer normalized output are
@@
where $\hat{x}^{(k)}=\frac{(x^{(k)}-\mu_{d})}{\sqrt{\sigma_{d}^{2}+\epsilon_{\sigma}}}$, $\mu_{d}=\frac{1}{d}\sum_{k=1}^{d}x^{(k)},\\ \sigma_{d}=\frac{1}{d}\sum_{k=1}^{d}(x^{(k)}-\mu_{d})^{2}$. Here, $\gamma^{(k)}$ and $\beta^{(k)}$ are learnable affine parameters (and we denote the set of them as $\lambda$), $\epsilon_{\sigma}$ is a small value to prevent numerical problems. Adaptation to new domains can be conducted by tuning $\gamma^{(k)}$ and $\beta^{(k)}$ in Eqn. 3.
4.4 Sharpness-aware domain adaptation
While TTA provides stability in LN models, it can lead to model collapse during affine parameter tuning (Eqn. 3). This collapse occurs when the model incorrectly assigns all input samples to a single class over the adaptation process[35]. To prevent this, we utilize sharpness-aware techniques [10, 35] to make the model less sensitive to large gradients and in test samples [34].
@@ 4.4 Sharpness-aware domain adaptation
We seek to make the model insensitive to the large gradients by encouraging the model to converge to a flat area of the entropy loss surface, since a flat minima has good generalization and robustness to large gradients[10, 35]:
@@
The abbreviation "SA" stands for sharpness-aware. In this context, the inner optimization aims to discover a perturbation $\mathbf{\epsilon}$ of LayerNorm affine parameters $\mathbf{\lambda}$ within a Euclidean ball of radius $\rho$ that maximizes entropy. The degree of sharpness is measured by the maximum change in the Euclidean ball neighbourhood $N_{\rho}(\mathbf{\lambda})$. This bi-level problem incentivizes the optimization process to locate flat minima. Following SAM[10], we can approximately solve the inner optimization via first-order Taylor expansion
@@
$\mathbf{\epsilon^{*}}(\mathbf{\lambda})=\rho\text{sgn}(\mathbf{v})\frac{|\mathbf{v}|}{\|\mathbf{v}\|_{2}}$ (7)
@@
By applying Eqn. 8 to update the LN affine parameters of $j$-th model with Eqn. 4, we can obtain a more reliable solution for entropy minimization. We update all the LayerNorm affine parameters of $j$-th source model by back propagating $\mathcal{L}_{j}^{(t)}$ once, as the approach in [52]. After updating $\gamma$s and $\beta$s of LN layers of $j$-th model, we denote the updated model as $\overline{f_{\theta_{j}}^{j(t)}}$ which will be used of the inference on $(t+1)$-th target test batch. This implies $f_{\theta_{j}}^{j(t+1)}=\overline{f_{\theta_{j}}^{j(t)}}$.
5 Experiment
Table 2: Results on Office-Home. We train the source models using 3 datasets to perform inference on the remaining dataset. The table clearly shows that PLUTO outperforms all of the baselines consistently
@@ 5 Experiment
Single Source-Best 62.5 47.7 74.3 69.4 63.4
Source-Worst 42.5 40.7 58.2 60.1 50.4
TENT-Best 65.5 51.2 76.8 71.4 66.2

### 08 Feb 2024 v3

@@ 3.2 Module weighting at test time
Drawing from the success of the attention mechanism in achieving impressive few-shot performance in supervised model ensemble scenarios (Peng et al., 2022; Kim et al., 2017), we design a novel approach to leverage the same during test time adaptation. Even in the unsupervised context of test-time adaptation, we have found that an attention module, which acts as a module selector, can effectively integrate information from multiple source domains into the target domain by dynamically assigning weights and updating them on-the-fly. This enables PLUTO to plug modules from the module store and adapt them at test time with few-shot unlabeled data.
@@
In order to convert a 2D image input into sequential data for the transformer and the module selector to process, the image input $x\in\mathbb{R}^{H\times W\times C}$ is reshaped into a sequence of flattened 2D patches (Dosovitskiy et al., 2020) $\mathbf{X}_{p}\in\mathbb{R}^{e\times(P^{2}\cdot C)}$, where $(H,W)$ is the height and width of the original image, $C$ is the number of channels, $(P,P)$ is the size of each segmented image patch, and $e=HW/P^{2}$ is number of patches and the effective input sequence length. Assuming that the transformer uses a constant latent vector size $d$ across all of its layers(Dosovitskiy et al., 2020), we flatten the sequence of the patches and map them to $d$ dimentions: $\hat{\mathbf{X}}=[\mathbf{X}_{p}\mathbf{E}]_{p=1}^{N},\mathbf{E}\in\mathbb{R}^{(P^{2}\cdot C)\times d}$. The patch embedding projection $\mathbf{E}$ is a (frozen) trainable parameter in the pretrained ViT.
@@
The module selector, which is an attention module, gauges the relevance between the output logits of the source modules and instances $\bm{x}_{i}$ in the current ($t$-th) target (abbreviated as $T$) batch $D_{T}^{(t)}=\{x_{i}^{(t)}\}_{i=1}^{B}$. It then assigns weights based on this relevance. Logits of source models are a function of both the pretrained module and the pretrained transformer and are also capable of modeling the uncertainty in predictions from different sources. We perform max-pooling over the sequence $\hat{\mathbf{X}}=[\mathbf{X}_{p}\mathbf{E}]_{p=1}^{N}\in\mathbb{R}^{e\times d}$ and obtain $\hat{x}\in\mathbb{R}^{d}$ as the representation of the original image input. Furthermore, we pass the input instance $x$ through $N$ pretrained models and obtain pre-softmax logits $\{l(x_{i})^{j}\}_{j=1}^{N}\in\mathbb{R}^{v}$. With four trainable parameters: $W_{d,x}\in\mathbb{R}^{d\times d^{\prime}_{x}},W_{u,x}\in\mathbb{R}^{d^{\prime}_{x}\times d^{\prime}},W_{d,l}\in\mathbb{R}^{v\times d^{\prime}_{l}},W_{u,l}\in\mathbb{R}^{d^{\prime}_{l}\times d^{\prime}}$, the module selector first projects the pre-softmax logits and the representation of the original image input into another representational space: $\mathbf{h}_{x}=W_{u,x}^{\intercal}\cdot\gamma(W_{d,x}^{\intercal}\cdot\hat{x}),\mathbf{h}_{l,j}=W_{u,l}^{\intercal}\cdot\gamma(W_{d,l}^{\intercal}\cdot{l(x)^{j}})$. With application of LN (Ba et al., 2016) we are able to obtain the final projected representations $\mathbf{h}_{x},\mathbf{h}_{l,j}\in\mathbf{R}^{d^{\prime}}$ and compute the attention weight $w(x)^{j}$ to capture the relationship between the input $x$ and its representation from the pretrained models $l(x)^{j}$:
@@
The ensembled logit is the linear combination of the weighted logits: $l(x)=\sum_{j=1}^{N}w(x)^{j}l(x)^{j}.$
As the GT label of the target sample $x$ is unknown, we denote its pseudo-label, as predicted by source $j$, as $\hat{y}_{j}=f^{j}_{\theta_{j}}(x)$. We linearly combine these pseudo-labels by attention weights $\{w(x)^{j}\}_{j=1}^{N}$ to get the weighted pseudo-label $\hat{y}=\sum_{j=1}^{N}w(x)^{j}\hat{y}_{j}$. Using these weighted pseudo-labels for all the samples in the current ($t$-th) batch, we calculate the expected Shannon entropy of pseudo-label as:
@@
where $K$ is the number of classes and $\hat{y}_{c}$ is the $c$-th entry of the predicted pseudo-label. By performing entropy minimization on Eqn. 2, we can update the module selector parameters (from $\mathcal{G}$ to $\mathcal{G}^{*}$) and obtain updated attention weights for Eqn. 1. Minimization of the Shannon entropy attempts to raise the confidence of an individual sample’s prediction and weight assignment (Lee et al., 2023).
@@ 3.3 Domain adaptation of LN affine parameters
In TTA, prior methods often conduct adaptation on pre-trained models with batch normalization (BN) layers (Ioffe & Szegedy, 2015) and most of them are built upon BN statistics adaptation (Schneider et al., 2020; Wang et al., 2020; Zhang et al., 2022). However, recent works (Liu et al., 2022; De Min et al., 2023; Kim et al., 2021) found out that layer normalization (LN, (Ba et al., 2016)) can be a better adaptation choice for transformers. Specifically, for the $k$-th layer with a $d$-dimensional input $x=(x^{(1)},\cdots,x^{(d)})^{\intercal}\in\mathbb{R}^{d}$, the layer normalized outputs are
@@
where $\hat{x}^{(k)}=\frac{(x^{(k)}-\mu_{d})}{\sqrt{\sigma_{d}^{2}+\epsilon_{\sigma}}}$, $\mu_{d}=\frac{1}{d}\sum_{k=1}^{d}x^{(k)},\\ \sigma_{d}=\frac{1}{d}\sum_{k=1}^{d}(x^{(k)}-\mu_{d})^{2}$. Here, $\gamma^{(k)}$ and $\beta^{(k)}$ are learnable affine parameters (and we denote them together as $\lambda$), $\epsilon_{\sigma}$ is a small value to ensure numerical convergence. Adaptation to new domains can be achieved by tuning $\gamma^{(k)}$ and $\beta^{(k)}$ in Eqn. 3.
3.3.1 Sharpness-aware domain adaptation
While TTA provides stability in LN models, it can lead to model collapse during affine parameter tuning (Eqn. 3). This collapse occurs when the model incorrectly assigns all input samples to a single class over the adaptation process (Niu et al., 2023). To prevent this, we utilize sharpness-aware techniques (Foret et al., 2020; Niu et al., 2023) to make the model less sensitive to large gradients and in test samples (Niu et al., 2022).
@@ 3.3.1 Sharpness-aware domain adaptation
We seek to make the model insensitive to the large gradients by encouraging the model to converge to a flat area of the entropy loss surface, since a flat minima has good generalization and robustness to large gradients (Foret et al., 2020; Niu et al., 2023):
@@
The abbreviation "SA" stands for sharpness-aware. In this context, the inner optimization aims to discover a perturbation $\mathbf{\epsilon}$ of LayerNorm affine parameters $\mathbf{\lambda}$ within a Euclidean ball of radius $\rho$ that maximizes entropy. The degree of sharpness is measured by the maximum change in the Euclidean ball neighbourhood $N_{\rho}(\mathbf{\lambda})$. This bi-level problem incentivizes the optimization process to locate flat minima. Following SAM (Foret et al., 2020), we can approximately solve the inner optimization via a first-order Taylor expansion,
@@
$\mathbf{\epsilon^{*}}(\mathbf{\lambda})=\rho\cdot\text{sgn}(\mathbf{v})\cdot\frac{|\mathbf{v}|}{\|\mathbf{v}\|_{2}}$ (7)
@@
By applying Eqn. 8 to update the LN affine parameters of $j$-th model with Eqn. 4, we can obtain a more reliable solution for entropy minimization. We update the LayerNorm affine parameters for all the layers of $j$-th source model by back propagating $\mathcal{L}_{j}^{(t)}$ once, as the approach in (Wang et al., 2020). After updating $\gamma$s and $\beta$s of LN layers of $j$-th model, we denote the updated model as $\overline{f_{\theta_{j}}^{j(t)}}$ which will be used of the inference on $(t+1)$-th target test batch. This implies $f_{\theta_{j}}^{j(t+1)}=\overline{f_{\theta_{j}}^{j(t)}}$.
4 Evaluations
Table 2: Results on Digits. We train the source modules using 4 digits datasets to perform inference on the remaining dataset ($M=N=4$). The table clearly shows that the average accuracy of PLUTO outperforms all of the baselines consistently.
Source Method MM MT UP SV SY Avg
Single TENT-Best 56.1 98.4 84.9 87.0 95.2 84.3
TENT-Worst 17.6 54.2 59.6 11.4 15.5 31.7
CoTTA-Best 55.1 97.5 84.9 86.3 93.6 83.5
CoTTA-Worst 17.2 54.2 59.7 10.8 15.6 31.5
EaTA-Best 54.8 97.3 84.2 85.4 93.0 82.9
EaTA-Worst 16.9 53.4 58.9 10.4 15.8 31.1
Multi TENT-Ens 62.4 97.8 87.6 55.1 78.4 76.3
CoTTA-Ens 55.7 97.9 85.7 87.1 95.5 84.4
EaTA-Ens 55.6 98.0 85.2 85.9 94.9 83.9
PLUTO (Ours) 63.1 98.9 89.6 87.4 93.7 86.6
Table 3: Results on Office-Home. We train the source modules using 3 datasets to perform inference on the remaining dataset ($M=N=3$). The table clearly shows that the average accuracy of PLUTO outperforms all of the baselines consistently.
@@ 4 Evaluations
Single TENT-Best 65.5 51.2 76.8 71.4 66.2
@@
CoTTA-Best 64.4 50.2 76.7 70.5 65.5
CoTTA-Worst 46.6 46.0 56.0 63.9 53.1
EaTA-Best 64.8 50.9 77.1 71.4 66.1
EaTA-Worst 47.4 46.7 55.4 62.8 53.1
@@
Table 4: Results on Digits. Different number of few-shot target samples for training the module selector and tuning the LayerNorm affine parameters.
Data Size Method MM MT UP SV SY Avg
128 TENT-Ens 62.4 97.8 87.6 55.1 78.4 76.3
CoTTA-Ens 55.7 97.9 85.7 87.1 95.5 84.4
EaTA-Ens 55.6 98.0 85.2 85.9 94.9 83.9
PLUTO 63.1 98.9 89.6 86.4 93.7 86.4
64 TENT-Ens 53.8 89.5 79.2 43.3 67.4 66.6
CoTTA-Ens 45.1 89.9 77.6 75.3 87.3 75.0
EaTA-Ens 43.7 87.1 73.5 76.0 86.6 73.4
PLUTO 59.8 94.6 86.5 81.8 88.3 82.2
16 TENT-Ens 47.8 84.2 73.0 37.6 60.5 60.6
CoTTA-Ens 38.3 82.4 71.5 68.8 82.0 68.6
EaTA-Ens 38 81.1 67.7 70.3 81.6 67.7
PLUTO 55.6 90.2 81.0 77.0 83.1 77.4
Zero-shot TENT-Ens 41.5 78.1 66.6 29.4 53.9 53.9
CoTTA-Ens 28.6 76.3 64.8 60.9 73.3 60.8
EaTA-Ens 28.6 71.9 58.5 61.8 72.9 58.8
PLUTO 51.5 84.6 73.8 71.8 76.6 71.7
Table 5: Results on Office-Home. Different number of few-shot target samples for training the module selector and tuning the LayerNorm affine parameters.
Data Size Method Ar Cl Pr Re Avg
128 TENT-Ens 63.1 49.2 75.6 71.0 64.7
CoTTA-Ens 64.6 50.4 76.4 71.7 65.8
EaTA-Ens 65.8 51.2 76.7 72.6 66.6
PLUTO 66.0 53.6 79.2 76.3 68.7
64 TENT-Ens 49.9 33.9 62.9 54.0 50.2
CoTTA-Ens 49.6 34.2 61.0 58.8 50.9
EaTA-Ens 51.4 35.9 62.8 58.7 52.2
PLUTO 59.7 47.5 71.7 70.2 62.3
16 TENT-Ens 31.5 15.0 47.6 37.4 32.9
CoTTA-Ens 32.4 17.7 42.2 43.6 34.0
EaTA-Ens 34.9 17.7 44.2 40.7 34.4
PLUTO 55.9 44.3 67.8 66.4 58.6
Zero-shot TENT-Ens 27.8 11.4 44.3 34.3 29.5
CoTTA-Ens 29.0 13.9 38.8 39.8 30.4
EaTA-Ens 30.9 13.7 41.0 37.2 30.7
PLUTO 52.4 40.9 63.9 62.4 54.9
Table 6: Results CIFAR-10C. We take $M=N=4$ source modules trained on Snow, Frost, Fog, and Bright. We employ these models for adaptation on 15 sequential target domains (severity level=1).
GN SN IN DB FGB MB ZB Snow Frost Fog Bright Contrast Elastic Pixel JPEG Avg
CoTTA-Best 58.0 59.3 43.9 54.6 45.7 43.0 43.8 37.0 36.1 27.5 29.0 14.3 14.4 12.9 9.7 35.3
EaTA-Best 55.2 57.6 45.5 66.7 58.5 66.3 69.3 69.7 72.9 70.6 69.2 67.4 62.6 68.1 52.4 63.5
TENT-Best 52.4 55.6 46.0 70.1 61.4 68 68.5 68.2 70.9 71.0 71.5 68.7 67.8 69.5 61.9 64.8
CoTTA-Ens 58.4 61.3 42.5 57.0 45.1 44.5 45 38.5 36.0 26.7 27.4 15.4 14.9 12.8 8.2 35.6
EaTA-Ens 56.2 58.5 46.7 69.2 56.6 68.3 70.4 69.2 75.5 71.5 68.9 69.9 64.7 67.1 53.8 64.4
TENT-Ens 53.6 57.7 45.8 71.9 59.5 68.4 68.8 69.5 72.2 70.5 69.7 69.7 67.8 69.7 62.4 65.1
PLUTO 61.6 63.3 48.7 75.2 64.9 71.8 73.7 73.0 78.9 75.0 74.9 73.2 70.5 72.8 65.0 69.5
Table 7: Results on ImageNet-C. PLUTO adaptation with top $k$ favorable source modules. $k$ stands for the number of used PET source modules in PLUTO, w.r.t the assigned weight. We adapt 15 target domains (severity level=1) separately.
Method k GN SN IN DB GB MB ZB Snow Frost Fog Bright Contrast Elastic Pixel JPEG Avg
PLUTO 1 66.2 75.4 76 74.9 74.9 77.2 75.9 75.2 76.6 76.4 76.3 76.3 75.1 77.2 76.8 75.4
3 69.8 78.3 78.7 77.6 76.9 79.5 77.9 77.6 79.3 79.2 78.6 79.3 77.7 79.4 79.3 77.9
5 72.3 78.9 79.2 78.3 77.9 80.3 78.2 78.1 80.4 80.5 79.2 80.0 78.1 79.9 79.7 78.7
15 72.7 79.3 79.7 78.7 78.3 80.9 78.8 78.8 81.0 80.8 79.8 80.5 78.6 80.4 80.4 79.2
75 72.8 79.5 80.1 78.9 78.5 81.1 79.1 79.0 81.3 81.0 80.0 80.7 78.8 80.5 80.5 79.4
In the main paper, we employ parameter-efficient fine-tuning (PET) in the form of VPT (Jia et al., 2022; Gao et al., 2022) as our module to demonstrate the effectiveness of our approach. We showcase our method’s application to other PET modules in the Appendix Sec. C to illustrate the generality of our approach. The effectiveness of our approach is evaluated on both stationary and continuous test time adaptation benchmarks. During the testing phase, we evaluate our approach under two scenarios: (1) Test batches are sampled from a stationary distribution (2) Test batches are sampled from a evolving dynamic distribution.
  • Digits-Five. (Peng et al., 2019) consists of five-digit datasets: MNIST (MT), MNIST-M (MM), USPS (UP), SVHN (SV), and Synthetic Digits (SY). There are 10 classes corresponding to digits ranging from 0 to 9. We employed four of the domains as sources and reserved the remaining domain for testing.
  • Office-Home. (Venkateswara et al., 2017) The Office-Home dataset consists of four domains: Art (Ar), Clipart (Cl), Product (Pr), and Real World (Re), each containing 65 classes. For evaluating our method in the test-time adaptation scenario, we selected three of these domains as source domains and used the remaining domain as the test data.
  • ImageNet-C. (Hendrycks & Dietterich, 2019) ImageNet-C dataset consists of 15 diverse corruption types applied to images of ImageNet (Deng et al., 2009). It consists of algorithmically generated corruptions with 5 severity levels. We regard different corruptions with different severity level as different domains. Thus we will have $5\times 15=75$ domains at most.
In the context of dynamic test distribution, we employ the following benchmarks.
  • CIFAR-100C. (Hendrycks & Dietterich, 2019) The CIFAR-100 (Krizhevsky et al., 2009) dataset is an image classification dataset consisting of 50,000 training images and 10,000 test images. Based on CIFAR-100, CIFAR-100C (Hendrycks & Dietterich, 2019) introduced 15 different types of noise at different severity levels (1 $\rightarrow$ 5). This dataset has since been utilized in various studies on continuous test-time adaptation.
4.1 Baseline Methods
Our evaluation includes comparisons with widely recognized single-source test-time adaptation methods, notably TENT (Wang et al., 2020), CoTTA (Wang et al., 2022), and EATA (Niu et al., 2022). These methods represent the current state-of-the-art (SOTA) in this area. We adopt a setup similar to that in (Ahmed et al., 2021), applying each source model individually to specific test domain data. This approach generates Best and Worst outcomes, corresponding to the highest and lowest performance achieved among the source models when adapted using a certain method, respectively. We also generalize these single-source adaptation methods into multi-source by uniformly averaging pre-softmax logits of all sources’ outputs given a target sample to make the prediction. The results of uniformly averaged Ensemble of multiple single sources methods are reported.
4.2 Implementation details
We use ViT-Base-16 (Dosovitskiy et al., 2020) model pretrained on JFT-300M dataset for all our experiments. For all experiments, we use a target batch size of $B=128$, as used by TENT (Wang et al., 2020). Further details on the experimental settings are provided in Appendix D.
4.3 Object recognition and digit classification
We report the results of object recognition on Office-home (Venkateswara et al., 2017) and Digits-Five (Peng et al., 2019) datasets in Table. 2,  3. We compute the accuracy for each incoming test batch and subsequently present the results by averaging the accuracy values across all the batches. For Office-home dataset, the number of source modules is $N=3$, and we use all $M=N$ (in Alg. 1) to do the inference on the target domain. A similar setup was used for Digits-Five dataset, except $M=N=4$ here.
4.4 Few-shot adaptation
Our default setting follows existing studies(Wang et al., 2020) by using 128 few-shot labeled data for each target task, as per the target batch size ($B$), to update the LN affine parameters and the module selector. Our interest lies in examining how PLUTO’s performance changes with varying shots ($U\leq B$) compared to other techniques. The outcomes are displayed in Table. 4 and Table. 5. We observe that PLUTO consistently outperforms existing TTA methods’ uniform ensemble across various amounts of few-shot unlabeled target data. It suggests that PLUTO can be applied to a wide range of few-shot settings. Even in zero-shot settings PLUTO can outperform other methods, since PLUTO benefits from its sample-specific adaptation strategy.
4.5 Forgetting issue
We extensively experiment with CIFAR-10C dataset to assess our model’s performance in the face of dynamic test distributions. Our method’s resistance to catastrophic forgetting is showcased by measuring classification accuracy on the original source test set after adapting to each domain (Niu et al., 2022; Song et al., 2023). For PLUTO, our ensemble method is employed during domain adaptation, and post-adaptation, each adapted source model is tested on its specific source test set. Regarding the baseline single-source methods and their average ensembles, each model is individually adapted to the new domain and then tested on its respective source test set. The reported accuracy in Table. 6 is the average derived from all these individually adapted single-source models. PLUTO serves an optimal aggregation of source models as well as better preservation of source knowledge.
4.6 Module selection
In previous experiments, we use all the source modules without selection for the target domains as mentioned earlier, regardless of their actual transferability given the target domain. However, if a source domain is too different from the target domain, it is possible that this source task wouldn’t effect the target task positively in general. In order to investigate how PLUTO performs when less preferable source domains are added, we conduct experiments as follows:
On ImageNet-C dataset, we pretrain a module store of $75$ modules($N=75$ in Alg. 1), each corresponding to a domain. We test PLUTO by choosing $M=k$ modules based on the module selector weight assignment for each module. The results presented in Table. 7, suggest that PLUTO effectively identifies target samples where less favored source tasks excel and appropriately adjusts sample-specific weights. This ability to rely more on these source models for certain target samples highlights the efficiency of PLUTO’s sample-specific approach.
5 Conclusion
PLUTO, our Test Time Adaptation (TTA) algorithm for transformers, adaptively combines several source parameter-efficient tuning (PET) modules during testing. Our experiments highlight PLUTO’s three key advantages: (1) Its adaptation strategy, which leverages multiple source domains, utilizes diverse domain knowledge for enhanced performance. (2) The sample-specific TTA approach of PLUTO contributes to its strong few-shot performance. (3) The use of PET modules for minimal parameter updates during test time prevents catastrophic forgetting. On multiple datasets, we demonstrate the power of our proposed approach, especially in a few-shot setting, and with a small number of selected modules.
Impact Statement
This paper presents a framework where a user can effectively adapt their machine learning model to their own needs by downloading modules from a global database. Thus, the outcomes of this work will potentially allow large transformer-based models to be applied to devices with limited computation power, e.g., those at the edge.
Appendix Overview
  • Section A Sharpness-aware domain adaptation
  • Section B Parameter efficiency
  • Section C Details on different PET methods
  • Section D Hyperparameters
  • Section E Details of dataset corruptions
  • Section F Theoretical insights
Appendix A Sharpness-aware domain adaptation
Following (Niu et al., 2023), in addition to minimizing the weighted pseudo-label entropy with Eqn. 8, we also seek to filter the samples that of large entropy value. As filtering samples based on their gradient norms directly is impractical, our approach focuses on examining the connection between entropy loss and gradient norms. We aim to eliminate samples with large gradients by considering their entropy values. The entropy is determined by the number of output classes, denoted as $K$, and it falls within the range $(0,\ln K)$ for various models and datasets. Consequently, selecting a threshold for filtering samples based on their entropy is a more manageable task. The threshold $E_{0}$ is set to $0.4\ln K$ by following EATA (Niu et al., 2022). $\rho$ in Eqn. 7 is set by the default value $5\times 10^{-2}$ in (Foret et al., 2020). We provide a overall view of the sharpness-aware DA applied in our method, PLUTO.
Algorithm 2 Sharpness-Aware Test-Time Entropy Minimization
1:Input: $t$-th target samples $D_{T}^{(t)}=\{x_{i}^{(t)}\}_{i=1}^{B}$, $j$-th source model $f^{j}_{\lambda}$ with tunable LN affine parameters $\lambda$, step size $\eta>0$, neighbourhood size $\rho>0$, entropy threshold $E_{0}$,
2:Output: updated LN parameter $\lambda^{*}$
3:for Each $x_{i}^{(t)}$ in $D_{T}^{(t)}$ do
4:     Predict pseudo-label $\hat{y}_{i}^{(t)}=f^{j}_{\lambda}(x_{i}^{(t)})$
5:     Compute entropy $\mathcal{E}_{i}^{(t)}$
6:      $\mathcal{E}_{i}^{(t)}(x_{i}^{(t)};\lambda)=-\sum_{c=1}^{K}\hat{y}_{ic}^{(t)}\log(\hat{y}_{ic}^{(t)})$
7:     if $\mathcal{E}_{i}^{(t)}>E_{0}$ then
8:         continue //filter the sample with large entropy
9:     end if
10:     Compute gradient $\mathbf{v}_{i}=\nabla_{\lambda}\mathcal{E}_{i}^{(t)}(x_{i}^{(t)};\lambda)$
11:     Compute $\mathbf{\epsilon}^{*}_{i}(\lambda)=\rho\text{sgn}(\mathbf{v}_{i})\frac{(\mathbf{v}_{i})}{\|(\mathbf{v}_{i})\|_{2}}$
12:     Compute gradient approximation:
13:     $\mathbf{g}_{i}=\nabla_{\lambda}\mathcal{E}_{i}^{(t)}(x_{i}^{(t)};\lambda)|_{\lambda+\mathbf{\epsilon}^{*}_{i}(\lambda)}$
14:     Update $\lambda\leftarrow\lambda-\eta\mathbf{g}_{i}$
15:end for
By applying Alg. 2, we minimize Eqn. 4 with Eqn. 8. On Office-Home dataset, we provide an ablation study of how much each component contribute to the performance in Table. A (Here, we apply VPT as the PET module, same as in the main paper Table. 3). We demonstrate the three components of PLUTO, namely, module selection, sharpness-aware pseudo-label entropy minimization (SA), and high-entropy sample filtering, through ablation experiments, highlighting their contributions to domain adaptation. It is evident that SA and Filter, in addition to module selection, contribute to the improved performance of PLUTO.